# Tracking electron capture processes in classical molecular dynamics simulations for spectral line broadening in plasmas

### v1

@@ IV.1 Application of the criterion
Let us now show an example of the application of this method, for the particular case of ions with charge $Z=2$, which means that the three closest neighbors must be monitored. In this example, the coupling parameter of the plasma is $\rho=0.6$, and therefore $\tau_{T}\sim 0.75$ in simulation units. Figure 1 shows the time-history of the potential energy of an emitter (top), the indices $I_{p}$ of the three electrons that are closest to it at every timestep in the simulation (middle) and the value of the electric field acting on said emitter (bottom panel). At $t\sim 2t_{0}$ (with $t_{0}$ being the time unit in the simulation), when the potential energy becomes lower than $\mathscr{E}_{th}$, the electron that is closest to the emitting ion is electron number 298. We then observe that this electron has been among the three closest neighbors from $t\sim 0$ to $t\sim 10t_{0}$, approximately 10 times the characteristic time of the simulation. This is much longer than the typical duration of a collision, so we must assume that this corresponds to a bound electron. Indeed, if we now consider the electric field between $t\sim 0$ and $t\sim 10t_{0}$, it oscillates periodically - with a period of approximately $0.75t_{0}$, which is the value of $\tau_{T}$ in this case - with very high values. A similar situation is observed at $t\sim 35t_{0}$, although in this case the trapped electron oscillates around the ion only a few times, and therefore its imprint on the electric field history is not as clear.
@@
It can be seen that this system does not eliminate strong collisions, as discussed above. For example, at $t\sim 20t_{0}$, the potential energy jumps below the threshold value, as an electron passes very close to the emitting ion. However, the colliding electron does not remain among the three closest neighbors for long enough to be a bound particle. Indeed, the electric field history confirms that this corresponds to an isolated collision.
@@
IV.2 Validation of the trapped-electron detection algorithm: calculation of the ionization balance
@@
Figure 2: Distribution of the potential energy of the ions (left) and its integral (right). Data correspond to a Helium plasma ($Z=2$), with an ionization potential $V_{i}=30\,\mathrm{\mathscr{E}_{0}}$, and a nominal coupling parameter $\rho_{N}=1.0$. The mean total energy per particle of the system is $\mathscr{E}_{t}=-2.12\mathscr{E}_{0}$.
@@
Figure 3: Same as in Fig. 2, with a mean total energy for particle of $-10.00\mathscr{E}_{0}$.
So far we have described the algorithm that allows tracking the electron trajectories in time and eventually determine whether a particular one has been trapped by an ion or not. Accordingly, this procedure allows identifying and storing the appropriate electric field sequences that will be further used for line shape calculations.
@@
It would be convenient to validate the trapped-electrons detection algorithm using a different independent method, also arisen and consistent with the FMD framework. In this regard, we note that by applying the trapped electron detection algorithm one can easily determine the fractional population of different ion charge states in the plasma at given conditions and therefore obtain the corresponding ionization balance.
It turns out that the plasma ionization balance can also be obtained from the analysis of the ion potential energy distribution. Once the equilibration time is achieved for a particular simulation run, one can now take a sufficiently large number of simulation snapshots during the equilibrium stage, compute the potential energy distribution for each of them and finally take the average in order to get a good statistical representation (smooth) of the potential energy distribution of the plasma particles at given conditions. To ensure proper statistics, additionally we take the average over a collection of simulation runs ($\sim 10$) starting with different initial conditions always consistent with the plasma conditions of interest. As explained below, the plasma ionization balance can be inferred by properly interpreting the potential energy distribution. We note that this is a robust method that does not require the definition of ad-hoc criteria to label an electron as bound or free. On the contrary, due to its own statistical nature, this method is not useful for selection of electric field sequences since the time-resolved information is lost. Hence, the comparison of the ionization balance obtained from the trapped-electron detection algorithm against the one from the analysis of the potential energy distribution provides a way to validate the former, adding confidence on the selected electric field sequences for line shape calculations.
@@
For illustration let us first assume the simple case of hydrogen ions ($Z=1$). If the plasma temperature is low enough (in simulation units) so that the distribution of kinetic energies is relatively narrow, the distribution of potential energies presents two main lobes. One of them centered around zero (free particles), and one of them near the potential well of energy $V_{i}$ (ions with a bound electron) - see figure 14 in Gigosos et al. [20] -. Following this logic, it can be easily seen how for ions with $Z=2$, the distribution of potential energies can have up to three lobes: that corresponding to free particles, around zero potential energy; the one accounting for ions with one bound electron, with potential energies similar to $V_{i}$; and the lobe corresponding to ions with two trapped electrons, which will have energies similar to $2V_{i}$. If these three groups are sufficiently separated, it is possible to integrate the energy distribution and associate the area under each lobe to the population of the different species.
@@
This is illustrated in Figure 2, in which we show the distribution of potential energies (left) and its integral (right) for an equilibrated plasma of $Z=2$. The corresponding potential well is $V_{i}=-30\mathscr{E}_{0}$, and the mean total energy per particle is $-2.12\mathscr{E}_{0}$. The distribution of potential energies therefore only presents two lobes, centered around zero and $V_{i}$. In this case, the energy per particle is relatively high compared with the ionization potential, and therefore only two ionization species are present in the plasma: free ions and those with one bound electron. By looking at the integral of the energy distribution (plot on the right), the ratio of ions with one bound electron ($I_{1}$) to free particles ($I_{0}$) can be easily obtained.
A similar example is shown in Figure 3 for the case when the plasma mean energy is $-10\mathscr{E}_{0}$. The rest of the plasma parameters are identical as those in Fig. 2. In this case, the plasma energy is too low, and the plasma is composed of only ions with one ($I_{1})$ or two bound electrons ($I_{2}$), with almost no free ions.
@@
Using this method, we calculated the fractional population distribution for a collection of simulation runs spanning over a wide range of plasma temperatures. The results are shown in Figure 4, where the the different colors correspond to the population fraction of the different ion species as a function of temperature. We note that each dot in fact corresponds to a group of different simulations associated to the same temperature value, i.e. for a given temperature, a number of simulations starting from different initial conditions were launched and then let to thermalize following the method described in our previous work [20]. As expected, at low temperatures all ions have two bound electrons and, as the temperature increases, some electrons become free, leading to an increase in the population of ions with only one bound electron. The number of ions with $Q=+1$ keeps growing until almost no neutral ions remain. If the temperature is increased further, ions can become doubly ionized ($Q=+2$), until there are no more bound electron-ion pairs.
Once the ion population distribution is calculated, one can easily obtain the average ionization. We show this in Figure 5, together with the results obtained from the application of the trapped-electron detection algorithm for the same temperature range. The agreement between the two methods is remarkable, which firmly constitutes a validity proof of the detection algorithm. Furthermore, as mentioned in the previous section, it can be seen that, even when the temperature becomes one tenth of the ionization potential, most atoms are fully ionized, and recombination effects become negligible.
V Conclusions
In this study, we have introduced a robust criterion for tracking and identifying trapped electrons within Full Molecular Dynamics (FMD), specifically focusing on ions with charges $Z>1$. Our presented criterion effectively distinguishes strong collisions within the plasma from recombination effects, providing a clear understanding of the underlying processes. Moreover, this method offers a pathway to determine the plasma ionization balance under specific conditions.
To validate our detection algorithm, we conducted extensive comparisons across a wide range of plasma temperatures. Specifically, we compared the average ionization values obtained through our criterion with those derived from a well-established method based on the analysis of ion potential energy distribution. It is important to note that while our ionization balance calculations do not claim to precisely solve plasma atomic kinetics, our full molecular dynamics simulations accurately capture the plasma ionization balance within the framework of the classical ionization-recombination mechanism. It is worth mentioning that this classical approach lacks a distinct quantum-mechanical counterpart.
@@ Acknowledgements
This work has also been carried out within the framework of the EUROfusion consortium, funded by the European Union via the Euratom Research and Training Program (Grant Agreement Nos. 633053 and 101052200—EUROfusion). The views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. The involved teams have operated within the framework of the Enabling Research Projects: Grant number AWP21-ENR-IFE.01.CEA.

### v2

@@ III.1 Application of the criterion
Let us now show an example of the application of this method, for the particular case of ions with charge $Z=2$, which means that the three closest neighbors must be monitored. In this example, the coupling parameter of the plasma is $\rho=0.6$, and therefore $\tau_{T}\sim 0.75t_{0}$ and $\tau_{\text{bound}}\sim 2.25t_{0}$. Figure 1 shows the time-history of the potential energy of an emitter (top), the indices $I_{p}$ of the three electrons that are closest to it at every timestep in the simulation (middle) and the value of the electric field acting on said emitter (bottom panel). The potential energy is given in units of the characteristic energy of the simulation $\mathscr{E}_{0}$, a parameter given by the average kinetic energy of the particles at the beginning of the simulation –before equilibrium is reached–, as further detailed in Gigosos et al. [15]. The electric field is given in units of $E_{0}$, i.e. the field produced by a single electron at the typical distance $r_{e}$.
At $t\sim 2.7t_{0}$, when the potential energy becomes lower than $\mathscr{E}_{th}$, the electron that is closest to the emitting ion is electron number 298. We then observe that this electron has been among the three closest neighbors from $t\sim 0$ to $t\sim 10t_{0}$, approximately 10 times the characteristic time of the simulation. This is much longer than the typical duration of a collision, so we must assume that this corresponds to a bound electron. Indeed, if we now consider the electric field between $t\sim 0$ and $t\sim 10t_{0}$, it oscillates periodically, with a period of approximately $0.75t_{0}$ (which is the value of $\tau_{T}$ in this case), reaching very high values. A similar situation is observed at $t\sim 40t_{0}$, although in this case the trapped electron oscillates around the ion only a few times, and therefore its imprint on the electric field history is not as clear.
It is clear now that the exact value of $\mathscr{E}_{th}$ is not critical, since strong collisions and electron captures appear in the time-history of the potential energy of the ion as clear spikes, and what determines whether an ion has captured an electron or not, is the time-dependent part of the criterion. In the example presented in Fig. 1, the ionization potential is $V_{i}=15\mathscr{E}_{0}$, which following the definition in Eqn. 3 corresponds to $\mathscr{E}_{th}=-10\mathscr{E}_{0}$. However, if its value were taken between $\sim-3\mathscr{E}_{0}$ and $\sim-12\mathscr{E}_{0}$, the same results would have bee obtained.
@@
In fact, by taking a lower threshold energy, the only difference is the computational cost of the simulation, since the time-history of the ion would need to be checked more often. For example, at $t\sim 20t_{0}$, the potential energy jumps below $6\mathscr{E}_{th}$, as an electron passes very close to the emitting ion. If the chosen threshold energy were lower, the algorithm would inspect this event as a potential electron capture. However, the colliding electron in this case does not remain among the three closest neighbors for long enough to be a bound particle. Indeed, the electric field history confirms that this corresponds to an isolated collision. It can therefore be seen that this system does not eliminate strong collisions and thus, by applying this criterion, it is possible to identify and select the appropriate sequences of the electric field that a radiator is subjected to during its emission, including the correlation loss owed to capturing an electron. These sequences can then be further used for Stark-broadening line shape calculations.
@@
III.2 Validation of the trapped-electron detection algorithm: calculation of the ionization balance
@@
Figure 2: Distribution of the potential energy of ions (left) and its integral (right). Data correspond to a Helium plasma ($Z=2$), with an ionization potential $V_{i}=30\,\mathrm{\mathscr{E}_{0}}$, and a coupling parameter $\rho_{N}=1.0$. The mean total energy per particle of the system is $\mathscr{E}_{t}=-2.12\mathscr{E}_{0}$.
@@
Figure 3: Same as in Fig. 2, with a mean total energy for particle of $10.00\mathscr{E}_{0}$.
@@
In order to validate the electron-trapping detection algorithm that we have established, we aim to compare its predictions with those from a different independent method, consistent with the FMD framework. In this regard, we note that this algorithm can additionally be used to determine the number of electrons that are trapped by an ion. As such, using this algorithm, one can easily determine the fractional population of different ion charge states in the plasma at given conditions and therefore obtain the corresponding ionization balance.
Nevertheless, the plasma ionization balance can also be obtained from the analysis of the ion potential energy distribution. In a FMD simulation with a given set of conditions, one can easily obtain the potential energy distribution of the plasma particles at any given time. By averaging over a sufficient number of instants and a sufficient number of simulation runs starting with different initial conditions ($\sim 10$), it is possible to obtain a good statistical representation of this distribution.
@@
To illustrate how the ionization balance can be obtained from the potential energy distribution, let us first assume the simple case of hydrogen ions ($Z=1$). If the plasma temperature is low enough so that the distribution of kinetic energies is relatively narrow, the distribution of potential energies presents two main lobes. One of them centered around zero (free particles), and one of them near the potential well of energy $V_{i}$ (ions with a bound electron) – as shown in figure 14 in Gigosos et al. [15] –. In the same manner, we could expect that, for ions with $Z=2$, the distribution of potential energies can have up to three lobes: that corresponding to free particles, around zero potential energy; the one accounting for ions with one bound electron, with potential energies similar to $V_{i}$; and the lobe corresponding to ions with two trapped electrons, which will have energies similar to $2V_{i}$ (and equivalently for higher $Z$).
If these lobes are sufficiently separated, it is possible to integrate the energy distribution and associate the area under each lobe to the population of the different species. This is a robust method that does not require the definition of ad-hoc criteria to label an electron as bound or free. On the contrary, due to its own statistical nature, this method is not useful for selection of electric field sequences since the time-resolved information is lost. Hence, the comparison of the ionization balance obtained from the trapped-electron detection algorithm against the one from the analysis of the potential energy distribution provides a way to validate the former, adding confidence on the selected electric field sequences for line shape calculations.
This is illustrated in Figure 2, in which we show the distribution of potential energies (left) and its integral (right) for an equilibrated plasma of $Z=2$. The corresponding ionization potential is $V_{i}=-30\mathscr{E}_{0}$, and the mean total energy per particle is $2.12\mathscr{E}_{0}$. In this case, the energy per particle is relatively high compared with the ionization potential, and therefore only two ionization species are present in the plasma: free ions and those with one bound electron. The distribution of potential energies therefore only presents two lobes, centered around zero and $V_{i}$. By looking at the integral of the energy distribution (plot on the right), the ratio of ions with one bound electron ($I_{1}$) to free particles ($I_{0}$) can be easily obtained.
A similar example is shown in Figure 3 for the case when the plasma mean energy is $10\mathscr{E}_{0}$. The rest of the plasma parameters are identical as those in Fig. 2. In this case, the plasma energy is too low, and the plasma is composed of only ions with one ($I_{1})$ or two bound electrons ($I_{2}$), with almost no free ions.
@@
Using this method, we calculated the fractional population distribution for a collection of simulation runs spanning over a wide range of plasma temperatures for the case of $Z=2$ and $V_{i}=-30\mathscr{E}_{0}$. The results are shown in Figure 4, where the the different colors correspond to the population fraction of the different ion species as a function of temperature. We note that each dot in fact corresponds to a group of different simulations associated to the same temperature value, i.e. for a given temperature, a number of simulations starting from different initial conditions were launched and then let to thermalize following the method described in our previous work [15]. As expected, at low temperatures all ions have two bound electrons and, as the temperature increases, some electrons become free, leading to an increase in the population of ions with only one bound electron. The number of ions with $Q=+1$ keeps growing until almost no neutral ions remain. If the temperature is increased further, ions can become doubly ionized ($Q=+2$), until there are no more bound electron-ion pairs.
Once the ion population distribution is calculated, one can easily obtain the average ionization, which permits the comparison with the results from the application of our trapped-electron detection algorithm. We show this in Figure 5 where the results from the two methods are shown for the same temperature range. The agreement between the two methods is remarkable, which firmly constitutes a validity proof of the electron capture detection algorithm. Additionally, as mentioned before, when the temperature becomes even one tenth of the ionization potential ($\sim 3\mathscr{E}_{0}$), it can be seen that most atoms are fully ionized, and recombination effects become negligible, which validates our estimation of $\tau_{bound}$.
IV Conclusions
@@
In this study, we have introduced a robust criterion for tracking and identifying trapped electrons in Full Molecular Dynamics (FMD) computer simulations, focusing on ions with charges $Z>1$. Our presented criterion effectively distinguishes strong collisions within the plasma from recombination effects, providing a clear understanding of the underlying processes, and allowing the selection of appropriate time-sequences of the electric field felt by a radiator for Stark broadening calculations. Additionally, this method offers a pathway to determine the plasma ionization balance under specific conditions.
To validate our detection algorithm, we conducted extensive comparisons across a wide range of plasma temperatures. Specifically, we compared the average ionization values obtained through our criterion with those derived from a well-established method based on the analysis of ion potential energy distribution. While our ionization balance calculations do not claim to precisely solve plasma atomic kinetics, our FMD simulations accurately capture the plasma ionization balance within the framework of the classical ionization-recombination mechanism. It is worth mentioning that this classical approach lacks a distinct quantum-mechanical counterpart.
@@
It is worth mentioning that the criterion presented in this work is applicable when the potential regularization radius $a$ is smaller than the average distance between particles $r_{e}$. While this is most often the case, for heavily coupled systems such as solid-density plasmas [19], care must be taken before enforcing such criterion. Future work on this topic will include the study of FMD simulations techniques for this type of crystalline systems.
@@ Acknowledgements
This work has also been carried out within the framework of the EUROfusion consortium, funded by the European Union via the Euratom Research and Training Program (Grant Agreement Nos. 633053 and 101052200—EUROfusion). The views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. The involved teams have operated within the framework of the Enabling Research Projects: Grant number AWP21-ENR-IFE.01.CEA. and AWP24-ENR-IFE.02.CEA-01.